Project

General

Profile

NDOS Commissioning Meeting 2012/02/20

Attendance: Jon, Minerba, Jan, Susan, Dave DeMuth

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the week
  3. DAQ Update and Plans for the week
  4. DCS Update and Plans for the week
  5. AOB

Beam and Detector Report

  • Peter: MI is currently down, but otherwise both beam and detector were within normal operating parameters this past week.

Installation plans for this week

  • No report

DAQ Update and Plans for next week

  • Peter: ~all software was updated on Friday. Frozen release defined. Problem with ups version of Pango, which is used by DAQ monitoring.
  • Jon: plans for this week? Peter: not really, perhaps some futher DDS tests, but those may also have been delayed.

DCS Update and Plans for next week

  • No report

AOB

  • Peter requested that a DataCheck update be given at these meetings. Jan has kindly agreed to do so.